twice the hundreds digit id the tens digit. the unit digit is four more than the hundreds digit. the sum of the number and the number formed by reversing the digits is 888. what is the numbertwice the hundreds digit id the tens digit. the unit digit is four more than the hundreds digit. the sum of the number and the number formed by reversing the digits is 888. what is the number
Answer by ankor@dixie-net.com(22740)

let a = the 100's digit
let b = the 10's digit
let c = the units
then
100a+10b+c = the number
:
Write an equation for each statement:
twice the hundreds digit is the tens digit.
2a = b
the unit digit is four more than the hundreds digit.
c = a + 4
the sum of the number and the number formed by reversing the digits is 888.
(100a + 10b + c) + (100c + 10b + a) = 888
combine like terms
100a + a + 10b + 10b + 100c + c = 888
101a + 20b + 101c = 888
replace b with 2a; replace c with (a+4)
101a + 20(2a) + 101(a+4) = 888
101a + 40a + 101a + 404 = 888
242a = 888 - 404
242a = 484
a = 484/242
a = 2
then
b = 2(2) = 4
and
c = 2 + 4 = 6
what is the number: 246 is the number
:
:
Check this 246 + 642 = 888
: